# Wave–Particle Complementarity as the Optimal Limit of Unambiguous Quantum-State Discrimination

### v2

@@ Abstract
We establish a direct operational interpretation of wave–particle complementarity in induced-coherence interferometry by linking interference visibility to the optimal performance of unambiguous quantum-state discrimination. In the low-gain Zou–Wang–Mandel interferometer, the idler modes of the two SPDC crystals act as nonorthogonal which-crystal marker states, and we show that the single-photon visibility equals the minimal inconclusive probability of the optimal Ivanovic–Dieks–Peres (IDP) strategy, $V=P_{I}^{\mathrm{opt}}$, so that the complementarity relation $D^{2}+V^{2}=1$ expresses a measurement-optimality boundary rather than a geometric constraint. Our results reveal that wave–particle duality in induced coherence is not limited by interferometer geometry but by measurement optimality in zero-error discrimination on the idler, with experimentally testable consequences predicting distinguishability beyond that inferred from visibility alone. In the presence of thermal noise we derive the hierarchy $V\leq F(\rho_{A},\rho_{B})\leq P_{I}^{\mathrm{opt}}$, linking visibility, fidelity, and optimal discrimination. These results identify distinguishability as fundamentally measurement-limited and motivate discrimination-enhanced induced-coherence imaging and sensing.
@@ I Introduction
Interferometry based on path indistinguishability, as realized in the seminal induced-coherence experiments of Zou, Wang, and Mandel (ZWM) [wang1991induced, zou1991induced, Zou1992], demonstrates that indistinguishable idler modes erase which-way information and induce first-order coherence between the corresponding signal fields. In the low-gain regime—where at most one of the two SPDC crystals emits a photon pair per pump pulse [Shafiee2024]—single-photon interference fringes appear in the signal arm despite the idler not being detected. Beyond fundamental interest, induced coherence is crucial in quantum imaging with undetected photons [Lemos2014], induced-coherence imaging theory [Lahiri2019], and phase-sensitive sensing [Shapiro2015].
The visibility $V$ of the interference fringes and the which-way distinguishability $D$ are linked by the complementarity relation $D^{2}+V^{2}=1$ for pure two-path markers [Wootters1979, Englert1996, Jaeger1993, mohseni], and recent measurements in induced-coherence systems confirm this behavior [Torres2024]. However, in all such treatments visibility is introduced as a geometric quantity associated with the coherence term of the reduced density matrix, without a direct operational interpretation in terms of a well-defined information-processing task.
In the low-gain ZWM interferometer, each photon pair originates from either crystal A or crystal B, but never both. This defines a natural binary hypothesis test [Helstrom]: Which crystal generated the pair? The corresponding conditional idler states are nonorthogonal, so perfect discrimination is impossible. Any zero-error strategy must therefore rely on unambiguous state discrimination (USD). For two pure states, the optimal USD strategy is the Ivanovic–Dieks–Peres (IDP) three-outcome POVM [Ivanovic, Dieks, Peres, Barnett].
Distinction from prior work.— Englert’s relation $D^{2}+V^{2}\leq 1$ expresses complementarity in terms of a geometric overlap between which-path marker states [Englert1996]. For pure marker states and ideal mode overlap, as realized in the low-gain ZWM interferometer [Torres2024], this bound is saturated, $D^{2}+V^{2}=1$. Our goal is to give this equality a direct operational interpretation by identifying the conventional visibility $V$ with the optimal inconclusive probability of the Ivanovic–Dieks–Peres (IDP) unambiguous discrimination measurement, $P_{I}^{\mathrm{opt}}$, on the conditional idler states. We further show that the ZWM geometry itself realizes the optimal three-outcome IDP POVM, without ancillary modes or adaptive stages, thereby connecting complementarity to a concrete decision-theoretic task. Bagan et al. [Bagan] formulated operational duality relations using discrimination games on abstract quantum states. In contrast, we work with a specific induced-coherence interferometer, identify the physically relevant which-crystal marker states, and show that the ZWM architecture provides a direct physical implementation of the optimal IDP POVM.
[…]
@@
Figure 1: Implementation of optimal unambiguous state discrimination (USD) [Barnett] in the induced-coherence (Zou–Wang–Mandel) interferometer [Torres2024]. Two coherently pumped nonlinear crystals (A and B) generate signal–idler photon pairs in the low-gain regime, so each emission event originates from either crystal A or crystal B but never both. The signal modes interfere at a balanced beam splitter $S_{1}$, producing single-photon fringes with visibility $V$. The idler modes ($\hat{a}_{3}^{2},\hat{a}_{3}^{3}$) form the two nonorthogonal marker states $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$, whose overlap determines $V$. The output of the first idler ($\hat{a}_{3}^{2}$) beam splitter is directed to detector $D_{0}$, realizing the inconclusive IDP outcome $\Pi_{0}$. The remaining idler amplitude is processed in a Mach–Zehnder network whose internal phases and beam-splitter ratios are tuned so that detectors $D_{1}$ and $D_{2}$ project onto the orthogonal complement states $|\phi_{2}^{\perp}\rangle$ and $|\phi_{1}^{\perp}\rangle$, yielding the conclusive IDP outcomes $\Pi_{1}$ and $\Pi_{2}$. This configuration realizes the optimal IDP POVM inside the nonlinear ZWM geometry and enables an experimental test of the operational identification $P_{I}^{\mathrm{opt}}=V$ and the complementarity relation $D^{2}+V^{2}=1$.
II ZWM Idler States and Hypothesis Testing
@@
ZWM marker states and complementarity.— In the low-gain regime [Shafiee2024, boyd, Kolobov2017], each pump pulse activates only one of the two nonlinear crystals, so each emission event corresponds to one of two mutually exclusive hypotheses. These hypotheses are encoded in the conditional idler states ($\hat{a}_{3}^{2},\hat{a}_{3}^{3}$) prepared by crystals A and B, which are generally nonorthogonal. As a result, perfect discrimination is impossible, and any zero-error identification must rely on an unambiguous state-discrimination (USD) strategy.
@@
$|\Psi\rangle\approx|0\rangle+g_{1}\,|1_{S_{1}},1_{I_{1}}\rangle+g_{2}\,|1_{S_{2}},1_{I_{2}}\rangle,$ (1)
with small pair-creation amplitudes $g_{1}$ and $g_{2}$. When the idler ($\hat{a}_{3}^{2},\hat{a}_{3}^{3}$) modes are aligned to be path-identical, $I_{1}$ and $I_{2}$ become a single physical mode $I$, and the biphoton component factorizes as
$\begin{split}g_{1}|1_{S_{1}},1_{I_{1}}\rangle+g_{2}|1_{S_{2}},1_{I_{2}}\rangle\;\longrightarrow\;\\[2.0pt] |1_{I}\rangle\otimes\big(g_{1}|1_{S_{1}},0_{S_{2}}\rangle+g_{2}|0_{S_{1}},1_{S_{2}}\rangle\big),\end{split}$ (2)
@@
We now introduce an object in the shared idler $I$ path between the two crystals, modeled as a lossless beam splitter with amplitude transmittance $t$ and reflectivity $r$, $|r|^{2}+|t|^{2}=1$. In the single-pair ($\hat{a}_{1}^{1},\hat{a}_{2}^{1}$) sector, the biphoton state at the output of the two crystals can be written (see Supplemental Material) as
$\begin{split}|\psi_{si}\rangle=\frac{1}{\sqrt{2}}\Big[&|1_{S_{1}},0_{S_{2}}\rangle\big(r|1_{I_{2}},0_{I_{3}}\rangle+t|0_{I_{2}},1_{I_{3}}\rangle\big)\\ &+e^{i\phi}|0_{S_{1}},1_{S_{2}}\rangle|0_{I_{2}},1_{I_{3}}\rangle\Big],\end{split}$ (3)
where $\phi$ is the relative pump phase. Conditioned on emission in crystal A (the $|1_{S_{1}},0_{S_{2}}\rangle$ branch), the idler state is
$|\phi_{1}\rangle=r\,|1_{I_{2}},0_{I_{3}}\rangle+t\,|0_{I_{2}},1_{I_{3}}\rangle,$ (4)
while emission in crystal B (the $|0_{S_{1}},1_{S_{2}}\rangle$ branch) leaves the idler in
$|\phi_{2}\rangle=|0_{I_{2}},1_{I_{3}}\rangle.$ (5)
These idler “marker” states carry which-crystal information; their overlap is
$\langle\phi_{1}|\phi_{2}\rangle=t.$ (6)
Tracing over the idler and recombining $S_{1}$ on a balanced signal beam splitter gives single-photon interference in the signal arm with visibility
$V=|\langle\phi_{1}|\phi_{2}\rangle|=|t|.$ (7)
@@
$D=\sqrt{1-|\langle\phi_{1}|\phi_{2}\rangle|^{2}}=|r|.$ (8)

### v3

# Which-crystal information and wave–particle duality in induced-coherence interferometry

@@ Abstract
We provide an operational reinterpretation of wave–particle complementarity in the low-gain Zou–Wang–Mandel (ZWM) induced-coherence interferometer. In the low gain limit, each photon pair is emitted by either one of two nonlinear crystals. Preparing nonorthogonal conditional idler states that encode which-crystal information. While previous studies inferred distinguishability indirectly from signal visibility with undetected idler photons. We show that the idler states naturally define a binary quantum hypothesis-testing problem. By performing optimal measurements on the idler, we analyze this task using both zero-error measurement unambiguous discrimination (Ivanovic–Dieks–Peres (IDP)) and minimum-error discrimination (Helstrom bound). We show that the signal visibility equals the optimal inconclusive probability of unambiguous discrimination. The Helstrom bound gives the optimal probability of identifying the emitting crystal. While signal visibility is an ensemble-averaged expectation value, the IDP and Helstrom strategies correspond to optimal single-photon decision measurements on the idler. The decision problem concerns inferring a past source event from a present measurement outcome. This establishes wave–particle duality in induced coherence as a manifestation of optimal quantum decision strategies rather than a purely geometric constraint. We further extend the analysis to the presence of thermal photons introduced in the object arm, which render the conditional idler states mixed. In this case, both the visibility and the achievable distinguishability are reduced, reflecting the fundamental limitations imposed by mixed-state discrimination. The approach is model-independent and applies to general two-path interferometers with markers.
@@ I Introduction
Interferometry based on path indistinguishability was introduced in the induced-coherence experiments of Zou, Wang, and Mandel (ZWM) [wang1991induced, zou1991induced, Zou1992]. In these experiments, indistinguishable idler modes erase which-way information and induce first-order coherence between spatially separated signal fields. Single-photon interference then appears in the signal arm, even when the idler photon is not detected [Shafiee2024]. This regime provides a natural platform for probing the quantum origin of induced coherence.
The quantum versus classical character of induced coherence has been the subject of active debate. Contextuality-based analyses identify measurement settings in which induced coherence leads to predictions incompatible with noncontextual hidden-variable models [Shafiee2024]. However, classical models reproduce several interference features of SPDC, including induced-coherence visibility in the low-gain regime [Kulkarni2022]. These results indicate that interference visibility alone does not fully characterize which-way or which-source information in induced-coherence interferometry.
The visibility $V_{Signal}$ of the signal interference fringes and the which-way distinguishability $D_{Signal}$ are linked by the complementarity relation $D^{2}+V^{2}=1$ for pure two-path markers [Wootters1979, Englert1996, Jaeger1993]. By contrast, Englert’s relation $D^{2}+V^{2}\leq 1$ formulates wave–particle complementarity in terms of the overlap of which-path marker states [Englert1996]. Bagan et al. [Bagan] further developed operational duality relations using discrimination games defined on abstract quantum states. However, the operational meaning of wave–particle complementarity in induced-coherence interferometry, particularly at the level of optimal single-photon inference, remains unclear.
A natural framework for addressing this limitation is provided by quantum decision theory. In particular, existing induced-coherence experiments do not access the full which-crystal information encoded in the idler state. In the low-gain ZWM interferometer, each photon pair originates from either crystal A or crystal B, but never both. This defines a binary quantum hypothesis test [Helstrom]: which crystal emitted the photon pair. The corresponding conditional idler states are generally nonorthogonal, so the previous discrimination is fundamentally constrained by quantum mechanics.
Optimal discrimination of nonorthogonal quantum states has been widely studied. Earlier experiments demonstrated the Ivanovic–Dieks–Peres (IDP) unambiguous discrimination measurement using optical interferometers [Clarke2001IDP, Herzog2005MixedUSD]. Related works have investigated path information and the past of photons inside interferometers using weak or frequency-based marking techniques [Zhou2017, Danan2013Trajectories], as well as unambiguous path discrimination in Mach–Zehnder interferometers [Len2018IDP]. These studies address questions of which-path information or spatial presence. In contrast, the present work focuses on induced-coherence interferometry and addresses a different retrodictive task: identifying which physical source emitted the photon pair.
In this work, we adopt a decision-theoretic perspective and analyze induced coherence using quantum state discrimination on the idler. For pure markers, we give the equality $D^{2}+V^{2}=1$ a direct operational meaning. We identify the fringe visibility $V_{\mathrm{Signal}}$ with the optimal inconclusive probability of the Ivanovic–Dieks–Peres (IDP) unambiguous discrimination measurement (USD), $P_{inc}^{\mathrm{opt}}$. We show that the ZWM geometry itself implements the optimal three-outcome IDP POVM [Ivanovic, Dieks, Peres, Barnett]. No ancillary modes or adaptive measurements are required. We also consider the complementary minimum-error strategy given by the Helstrom bound. This measurement maximizes the probability of correctly identifying which crystal emitted the photon pair. The resulting distinguishability quantifies which-crystal information in the minimum-error sense. Together, the IDP and Helstrom strategies yield complementary operational interpretations of wave–particle duality in induced coherence.
@@
Figure 1: Implementation of optimal unambiguous state discrimination (USD) [Barnett, Clarke2001IDP, Len2018IDP] in the induced-coherence (Zou–Wang–Mandel) interferometer [Torres2024]. Two coherently pumped nonlinear crystals (A and B) generate signal–idler photon pairs in the low-gain regime. Each emission event originates from either crystal A or crystal B but never both. The signal modes interfere at a balanced beam splitter $S_{1}$, producing single-photon fringes with visibility $V$. The idler modes ($\hat{a}_{3}^{2},\hat{a}_{3}^{3}$) form the two nonorthogonal marker states $|\phi_{A}\rangle$ and $|\phi_{B}\rangle$, whose overlap determines $V$. The output of the first idler ($\hat{a}_{3}^{2}$) beam splitter is directed to detector $D_{0}$, realizing the inconclusive IDP outcome $\Pi_{inc}$. The detectors $D_{1}$ and $D_{2}$ project onto the orthogonal complement states $|\phi_{B}^{\perp}\rangle$ and $|\phi_{A}^{\perp}\rangle$, yielding the conclusive IDP outcomes $\Pi_{A}$ and $\Pi_{B}$.
II Induced-coherence interferometer marker states and complementarity
@@
In the ZWM setup as shown in Fig. 1, coherently pump the two nonlinear crystals. In the low-gain regime [Shafiee2024, boyd, Kolobov2017, giri1], each emission event corresponds to one of two mutually exclusive hypotheses. These hypotheses are encoded in the conditional idler states prepared by crystals A and B, which are generally nonorthogonal. As a result, perfect discrimination is impossible. Any zero-error identification must rely on an unambiguous state-discrimination (USD) strategy.
@@
$|\Psi\rangle_{total}\approx|0\rangle+g_{A}\,|1_{S_{A}},1_{I_{A}}\rangle+g_{B}\,|1_{S_{B}},1_{I_{B}}\rangle,$ (1)
with small pair-creation amplitudes $g_{A}$ and $g_{B}$. When the idler modes ($\hat{a}_{3}^{1}=I_{A}$) are aligned to be path-identical, $\hat{a}_{3}^{2}=I_{A}^{{}^{\prime}}$ and $\hat{a}_{3}^{3}=I_{B}$ become a single physical mode $I_{A}^{{}^{\prime}}=I_{B}=I$, and the biphoton component factorizes as
$\begin{split}g_{A}|1_{S_{A}},1_{I_{A}}\rangle+g_{B}|1_{S_{B}},1_{I_{B}}\rangle\;\longrightarrow\;\\[2.0pt] |1_{I}\rangle\otimes\big(g_{A}|1_{S_{A}},0_{S_{B}}\rangle+g_{B}|0_{S_{A}},1_{S_{B}}\rangle\big),\end{split}$ (2)
@@
We now introduce an object in the shared idler $I$ path between the two crystals, modeled as a lossless beam splitter with amplitude transmittance $t$ and reflectivity $r$, $|r|^{2}+|t|^{2}=1$. In the single-pair sector, the biphoton state at the output of the two crystals can be written (see Appendix A) as
$\begin{split}|\psi\rangle_{SI}=\frac{1}{\sqrt{2}}\Big[&|1_{S_{A}},0_{S_{B}}\rangle\big(r|1_{I_{A}^{{}^{\prime}}},0_{I_{B}}\rangle+t|0_{I_{A}^{{}^{\prime}}},1_{I_{B}}\rangle\big)\\ &+e^{i\phi}|0_{S_{A}},1_{S_{B}}\rangle|0_{I_{A}^{{}^{\prime}}},1_{I_{B}}\rangle\Big],\end{split}$ (3)
where $\phi$ is the relative pump phase. Conditioned on emission in crystal A (the $|1_{S_{A}},0_{S_{B}}\rangle$ branch), the idler state is
$|\phi_{A}\rangle=r\,|1_{I_{A}^{{}^{\prime}}},0_{I_{B}}\rangle+t\,|0_{I_{A}^{{}^{\prime}}},1_{I_{B}}\rangle,$ (4)
while emission in crystal B (the $|0_{S_{A}},1_{S_{B}}\rangle$ branch) leaves the idler in
$|\phi_{B}\rangle=|0_{I_{A}^{{}^{\prime}}},1_{I_{B}}\rangle.$ (5)
Tracing over the idler and recombining $S_{1}$ on a balanced signal beam splitter gives single-photon interference in the signal arm $F$ with visibility
$V=|\langle\phi_{A}|\phi_{B}\rangle|=|t|.$ (6)
@@
$D=\sqrt{1-|\langle\phi_{A}|\phi_{B}\rangle|^{2}}=|r|.$ (7)
@@
An extension to imperfect idler mode overlap is given in Appendix B and leads to the reduction $V\rightarrow|\gamma|\,|t|$, with an unchanged complementarity relation $D^{2}+V^{2}=1$ in the pure-state model.
In Eq. (2), signal written as a coherent superposition of two emission alternatives. The path identical idlers $I$ are creating a two assumption, with object and without object.
If object $S$ is not present, idler modes ($\hat{a}_{3}^{3}=I_{A}$) associated with the two crystals are rendered path-identical $I$, the conditional idler states are identical, $|\phi_{A}\rangle=|\phi_{B}\rangle$. Tracing over the idler degree of freedom then yields a reduced signal state with a non-vanishing off-diagonal coherence term. Resulting in maximal first-order interference in the signal arm. This is the basic mechanism of induced coherence in the ZWM interferometer. On the other hand, the idler modes remain distinguishable and occupy orthogonal spatial modes. The conditional idler states are orthogonal, $\langle\phi_{A}|\phi_{B}\rangle=0$. In this case, tracing over the idler removes all coherence between the two signal emission alternatives, and no first-order interference is observed. Thus, path identity of the idler modes is a necessary condition for induced coherence.
If the object $S$ present, the common idler $I$ path modifies the idler associated with only one emission alternative. Although the idler modes remain path-identical. The object renders the conditional idler states nonorthogonal, $0<|\langle\phi_{A}|\phi_{B}\rangle|<1$. After tracing over the idler, the reduced density matrix of the signal contains an off-diagonal coherence term proportional to this overlap,
$\rho_{S_{A},S_{B}}\propto\langle\phi_{A}|\phi_{B}\rangle.$ (9)
As a result, the signal exhibits partial induced coherence with a visibility determined by the degree of nonorthogonality of the idler states.
If the idler modes $I$ are not path-identical. The conditional idler states remain orthogonal regardless of the presence of the object, and induced coherence is completely suppressed. Thus, path identity enables induced coherence, while the object controls its degradation by introducing controlled nonorthogonality between the conditional idler states.